OpenStudy (highschoolmom2010):

13. Using 6400 km as the radius of Earth, calculate how high above Earth’s surface you would have to be in order to weigh 1/16th of your current weight.

OpenStudy (anonymous):

You know this formula? \[F=\frac{ GMm }{ r^2 }\]

OpenStudy (highschoolmom2010):

i know of it but not too much on how to apply it

OpenStudy (anonymous):

ok. in G is a gravitation constant, M is the mass of the bigger body (Earth) and m is the mass of the smaller one (you). This application is comparing your weight on the Earth to somewhere in space so you can make a ratio. \[F _{earth}=\frac{ GMm }{ r^2_{earth} }\] \[F _{orbit}=\frac{ GMm }{ r^2 }\] \[\frac{ F _{orbit} }{ F _{earth} }=\frac{ 1 }{ 16 }\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):